OpenStudy (anonymous):

I have tried everything. I know the formula and how to solve it but no matter what my test keeps saying it is wrong. Find the volume of a cylinder with a diameter of 6 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank)

OpenStudy (ranga):

Volume of a cylinder = (pi)(r^2)h Diameter = 6. Radius is half the diameter. So radius r = 3 Height is three times the radius. So height h = 3 x r = 3 x 3 = 9 Just plug it into the formula for volume.

OpenStudy (anonymous):

I got 169.56

OpenStudy (ranga):

Volume = (pi)(r^2)h = (3.14)(3^2)(9) = 3.14 x 9 x 9 = 254.34

OpenStudy (anonymous):

Oh I see what I did wrong.I squared 3 wrong by mistake

OpenStudy (ranga):

But you have to square 3. radius r = 3 and you have to square it to get 9 as I did above.

OpenStudy (anonymous):

Right, okay I see. Thanks a lot!

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Would you mind helping me with another one?